i am having a major examine and trying to get ready for it but i am having some problems solveing the following type of problems can you show me
910cg = ? dag
7kg 25 g =? kg
5.311=?kl
2.4 sq m = ? sq cm

i appreciate your help very much

Answer by sarah_adam(201)

10 cg = 1 dg
910 cg = 910/10 = 91 dg
7kg 25 g =? kg
1000 g = 1kg
25 g = 25/1000 = 1/4 kg = 0.025kg
7kg 25 g = 7+0.025 = 7.025kg
5.311=?kl
1kl = 1000 l
5.3 l = 5.3/1000 = 0.0053 kl
2.4 sq m = ? sq cm
1 sqm = 100 sq cm
2.4 sq m = 2.4 *100 sq cm
2.4 sq m = 240 sq cm